# Negative non-Gaussianity as a salvager for PBHs with PTAs in bounce

### v1

@@ I Introduction
Recent advancements by the PTA experiments has marked a significant milestone with the detection of the Stochastic Gravitational Wave Background (SGWB). NANOGrav [1, 2, 3, 4, 5, 6, 7, 8], EPTA [9, 10, 11, 12, 13], CPTA [14], PPTA [15, 16, 17], have provided compelling evidence of the Hellings-Downs correlation pattern, strong signature of a gravitational wave origin. There are several possible sources to explain the origin of the SGWB. Some of the prominent ones include supermassive black holes, cosmic strings, domain walls, cosmic first order phase transitions, and induced Gravitational Waves (GW) from large quantum fluctuations, primarily the Scalar Induced Gravitational Waves (SIGW). For details on the above, please refer to [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 51, 40, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 33, 78, 79, 80, 81, 71, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121]. In this paper, we have focused on generating the SIGWs from the curvature perturbation. They do not interact with the surrounding plasma and, therefore, act as a crucial probe of the early universe (typically covering the scales $k\sim(10^{7}-10^{18}){\rm Mpc^{-1}}$). These SIGWs are tensor perturbations sourced at the second order in cosmological perturbation theory by the non-linear interactions between the scalar perturbations. Because the GWs come in the second order, the corresponding curvature perturbation needs to be significantly enhanced at small scales, compared to the Cosmic Microwave Background (CMB) scales, for possible detection by PTA experiments. This enhancement of the perturbation is also associated with formation of cosmic underdense and overdense patches within an otherwise isotropic and homogeneous FLRW background, which gravitationally collapse beyond a certain critical threshold that leads to the creation of Primordial Black Holes (PBHs)[122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 128, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 185, 186, 187, 202, 194, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 213, 207, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 232, 128, 233, 234, 235, 236, 237, 21, 22, 18, 19, 20, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 32, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 262, 263, 260, 264, 265, 266, 267, 268, 269, 29, 270, 271, 272, 81, 273, 274, 275, 276, 277, 278, 279]. PBHs come through as an intriguing alternative to explain a sizeable portion of the complete dark matter of the universe. [171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187]. However, there is a caveat that demands thorough scrutiny. Much like the most promising scenarios, the drawback of PBH formation is the risk of their overproduction [23, 280, 24, 281, 282, 82, 283, 71, 284, 285, 27, 286, 287, 18, 19, 20], a scenario where an abundance of PBHs, denoted by $f_{\rm PBH}$, appear to become more than unity; their energy density exceeds the totality of dark matter. Evading PBHs overproduction is a challenging task, more so in the context of SIGW interpretation of the PTA experimental data. Following are some references that have come up with a solution to this issue - [23, 280, 24, 281, 282, 82, 283, 71, 284, 285, 27, 286, 287, 18, 19, 20].
@@
PBHs are formed from large amplification of the curvature perturbation at specific scales. To this day, the exact scenario of PBHs formation is a largely discussed topic [288, 289]. The calculation of PBH abundance can be thought to be indirectly derived from knowledge of either of the two profiles: density perturbation profile and the curvature perturbation profile. Standard production mechanism involves the density perturbation amplitude averaged over a spherical region, crossing a certain threshold $\delta_{c}$, before collapsing into PBHs. In the traditional Press Schechter (PS) formalism [290], a Gaussian distribution of the density perturbation is assumed and is generally applicable to spherically symmetric models. In other words, the formalism is dependent on the shape of the perturbation profile. A physically more robust procedure to estimate the PBHs abundance has been developed known as the compaction function approach, which we here integrate in our study. The compaction function is similar to the schwarzschild gravitational potential, since it mainly relies on the functional form of the gravitational potential(or mass excess) within a region. In [291], it was shown that the threshold of the PBHs formation is largely unaffected beyond the critical radius $r_{p}$, denoting the peak of the compaction function. A crucial takeaway of this approach is that there is universal threshold value of the volume-averaged compaction function which can be easily related to $\delta_{c}$, and is independent of the perturbation profile to a high accuracy estimate [291], unlike in the PS method which relies on direct $\delta_{c}$ value and is heavily dependent on the perturbation shape. Additionally, PBH formation is a non-linear process and simple consideration of a Gaussian profile of density perturbation as in PS theory is unrealistic. In contrast, compaction function utilizes a non-linear relationship between the density and curvature perturbation in the long wavelength approximation which we discuss later in detail.
@@
Another important aspect of this work comes from considering the impact of quantum loop effects while studying PBH formation. More precisely, the scenario where huge enhancements in the curvature perturbation can induce significant quantum corrections at large scales whilst contributing to PBH formation was first investigated by Kristiano and Yokoyama in [292, 293]. This scenario quickly became a topic of intense discussion, followed by works such as [230, 231, 262, 261, 294, 295, 296, 297, 298, 299, 263, 260] discussing the proper conditions to conclude whether or not PBH formation remains ruled out. Parameters that include the duration of the perturbation amplifying phase, also known as the ultra-slow roll (USR) phase, and the nature (smooth or sharp) and implications of the transition in and out of this phase demand much scrutiny before establishing any firm results. In the midst of this, in a series of papers by the authors of [233, 235, 234], proposed a strong no-go theorem as a constraint on the largest mass of the PBHs, $M_{\rm PBH}\sim{\cal O}({10^{2}}\rm gm)$, as a result of the USR phase. The calculation mechanism involved the application of regularization, then renormalization, followed by a dynamical group resummation, to arrive at a finite result for the two-point correlation function of importance to us in the late time. Moreover, the underlying effective field theory (EFT) treatment of the no-go theorem propelled these conclusions to cover a wide range of canonical and non-canonical single-field inflationary models.
@@
Nevertheless, several studies like [236, 237, 21, 22, 19, 20, 239, 240] have demonstrated methods to circumvent this rigid PBH mass constraint using the EFT framework. Among such diverse set-ups, our paper focuses on a specific framework with additional pre-inflationary phases of contraction and an expanding non-singular bounce before following up with the standard SR-USR-SR stages for PBH generation. For our purposes, we utilize the result of performing the same calculation scheme as explained above to obtain the regularized-renormalized-resummed (RRR) scalar power spectrum, [240], within the EFT setting. The motivation here to introduce such phases in our set-up stems mainly from the fact that other than allowing us to evade the no-go, such phases have been shown to provide non-Gaussianities (NGs) that are orders of magnitude of larger than what is attainable in a standard single-field slow-roll inflation, as first shown from Maldacena’s calculations in [300] with $f_{\rm NL}\sim{\cal O}(0.1)$, or even, in scenarios with slow-roll violating phases like the USR with $101<|f_{\rm NL}|<$ 10. The matter bounce is such a scenario where the NGs are much larger (and negative) compared to standard inflationary scenarios, primarily because the curvature perturbation here grows on scales beyond the Hubble scale which later contributes to the strength and shape of the NGs [301]. Likewise, for models focusing on ekpyrotic contraction [302], negative NGs are also shown to be larger than the standard SR inflation. At this point, it might be tempting to ask that - with the introduction of all these phases, how large of a non-Gaussianity can be realistically feasible? Before answering that, let us briefly consider the importance of NGs in the context of PBH formation.
@@
The abundance of PBHs is affected by the Probability Distribution (PDF) of the curvature perturbation, which generally is assumed to be Gaussian. However, for a realistic analysis, the inclusion of NGs in the curvature perturbation, $\zeta$, is necessary to accurately estimate the abundance of PBHs, largely because there is no concrete confirmation that curvature perturbation itself behaves as a Gaussian random field. They emerge from quantum fluctuations and are responsible for creating a myriad of exotic and large-scale structures in the universe. Therefore, it is highly likely that the perturbation field would inherit intrinsic NGs. Besides, even under an ideal Gaussian assumption, the relationship between the curvature perturbation and the density contrast field is non-linear, which imparts NGs to the density contrast field as an effect due to these non-linear (NL) corrections. Previous case studies in [23, 20, 19, 18, 303] demonstrate that incorporating NGs have enabled a reduction in the abundance of PBHs, eliminating the risk of the overproduction. Most fiducial cases in the literature treat NGs perturbatively, while there have been some development in its study through non-perturbative formalism [304, 305, 306].
@@
In this paper, we will consider the local form of NGs in the curvature perturbation expressed with the parameter $f_{\rm NL}$ that encodes the first-order NG correction. PBH abundance is highly dependent on the signature of this $f_{\rm NL}$, with most of the cases displaying that $f_{\rm NL}<0$ is considered more favourable to suppress the PBHs abundance while positive NGs $f_{\rm NL}>0$ invites the risk of overproduction. In our case, such NGs are affected by the USR phase, usually short-lived with $\Delta N_{\rm USR}\sim 2$, and operate to enhance the fluctuations at small scales and violating Gaussianity. One could also conceive of a curvaton scenario [23, 307, 308, 306], where the curvature perturbation acquires NG corrections. Apart from this, scenarios described a while ago with ekpyrotic contraction and matter bounce also invites large amount of negative non-Gaussianity. The analysis in this paper incorporates the above mentioned two phases along with the three phases of the usual SR inflation model. Lets revert back to the question we asked a while ago about how large of a negative NG can we dive into. In [302], it has been shown through the calculation using the $\delta N$ formalism [309, 310, 311, 312, 313, 314, 315, 316, 317, 318, 319, 320, 321, 322, 323, 324, 325, 326, 304, 327, 328, 329], that $f_{\rm NL}$ can take values as large as $f_{\rm NL}\sim-166$ from the ekpyrotic phase. While in [301], it has been shown that $f_{\rm NL}$ is negative and of the order of $f_{\rm NL}\sim{\cal O}(1)$. When these phases along with an added USR phase is included, a large NG has to be expected. For our analysis, we adopt two benchmark values of $f_{\rm NL}\in(-39.95,-35/8)$ and present a comparative analysis to conclude with a stringent bound for the leading order NG parameter $f_{\rm NL}$. We will demonstrate through our analysis that diving below $f_{\rm NL}=-39.95$ for a specific threshold of the compaction function, raises the risk of breaking perturbativity conditions. We will confirm that with $f_{\rm NL}=-39.95$, a substantial amplification of the power spectrum is possible that explains the PTA and simultaneously avoid overproduction. Throughout this work, we have employed two distinct value of the effective sound speed parameter $c_{s}\in(0.88,1)$, lying within the window that preserves causality and unitarity conditions. The mass spectrum of PBHs examined in this work is between $M_{\rm PBH}\sim(10^{-6}-0.1)M_{\odot}$. Our set-up is successful in generating large mass PBHs while satisfying the inflationary conditions and underlying perturbativity conditions. We also clarify that in this work, the equation of state parameter has been fixed to a constant value corresponding to the radiation dominated epoch.
@@
This paper is structured as follows. We commence with a quick description of our EFT set-up in sec.(II). This is followed up with the expressions for the tree level and loop corrected scalar power spectrum of the curvature perturbation in sec.(III). In sec.(IV), the overall dynamics of PBHs formation has been discussed. Special emphasis has been put towards the development of the compaction function approach as it is going to be our prime tool throughout this paper. The formulas for the corresponding joint PDF and PBHs abundance has been laid down along with the domain of integration for the PBHs mass fraction. This is followed up with highlighting the main expressions for the SIGW spectrum in sec.(V) for $w=1/3$. We then present the outcomes of our numerical analysis in sec.(VI), where motivation for consideration of large negative NGs for the analysis has been largely discussed. To familiarize the reader with a rough idea of the recent studies, we have presented a light comparison in sec.(VII) between several models. Various experiments that our analysis has confronted have been enlisted in sec.(VIII). Based on our examination, we provide a stringent upper bound on the NG parameter $f_{\rm NL}$, with its validity, in sec.(IX). Finally, we draw our conclusion and highlight the major findings from this work in sec.(X).

### v3

@@ I Introduction
Recent advancements by the PTA experiments has marked a significant milestone with the detection of the Stochastic Gravitational Wave Background (SGWB). NANOGrav [1, 2, 3, 4, 5, 6, 7, 8], EPTA [9, 10, 11, 12, 13], CPTA [14], PPTA [15, 16, 17], have provided compelling evidence of the Hellings-Downs correlation pattern, strong signature of a gravitational wave origin. There are several possible sources to explain the origin of the SGWB. Some of the prominent ones include supermassive black holes, cosmic strings, domain walls, cosmic first order phase transitions, and induced Gravitational Waves (GW) from large quantum fluctuations, primarily the Scalar Induced Gravitational Waves (SIGW). For details on the above, please refer to [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 51, 40, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 33, 78, 79, 80, 81, 71, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 81, 128, 129, 130, 131, 132]. In this paper, we have focused on generating the SIGWs from the curvature perturbation. They do not interact with the surrounding plasma and, therefore, act as a crucial probe of the early universe (typically covering the scales $k\sim(10^{7}-10^{18}){\rm Mpc^{-1}}$). These SIGWs are tensor perturbations sourced at the second order in cosmological perturbation theory by the non-linear interactions between the scalar perturbations. Because the GWs come in the second order, the corresponding curvature perturbation needs to be significantly enhanced at small scales, compared to the Cosmic Microwave Background (CMB) scales, for possible detection by PTA experiments. This enhancement of the perturbation is also associated with formation of cosmic underdense and overdense patches within an otherwise isotropic and homogeneous FLRW background, which gravitationally collapse beyond a certain critical threshold that leads to the creation of Primordial Black Holes (PBHs)[133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 139, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 196, 197, 198, 213, 205, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 218, 225, 226, 227, 228, 229, 230, 231, 232, 233, 234, 235, 236, 237, 238, 239, 240, 241, 242, 243, 139, 244, 245, 246, 247, 248, 21, 22, 18, 19, 20, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 262, 263, 32, 264, 265, 266, 267, 268, 269, 270, 271, 272, 273, 274, 275, 276, 273, 277, 278, 279, 280, 281, 282, 29, 283, 284, 285, 81, 286, 287, 288, 289, 290, 291, 292, 293, 294, 295, 81, 128, 129, 130, 131, 132]. PBHs come through as an intriguing alternative to explain a sizeable portion of the complete dark matter of the universe [182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198]. However, there is a caveat that demands thorough scrutiny. Much like the most promising scenarios, the drawback of PBH formation is the risk of their overproduction [23, 296, 24, 297, 298, 82, 299, 71, 300, 301, 27, 302, 303, 18, 19, 20], a scenario where an abundance of PBHs, denoted by $f_{\rm PBH}$, appear to become more than unity; their energy density exceeds the totality of dark matter. Evading PBHs overproduction is a challenging task, more so in the context of SIGW interpretation of the PTA experimental data. Following are some references that have come up with a solution to this issue - [23, 296, 24, 297, 298, 82, 299, 71, 300, 301, 27, 302, 303, 18, 19, 20].
@@
PBHs are formed from large amplification of the curvature perturbation at specific scales. To this day, the exact scenario of PBH formation is a largely discussed topic [304, 305]. The calculation of PBH abundance can be thought to be indirectly derived from knowledge of either of the two profiles: the density perturbation profile and the curvature perturbation profile. The standard production mechanism involves the density perturbation amplitude averaged over a spherical region, crossing a certain threshold $\delta_{c}$, before collapsing into PBHs. In the traditional Press-Schechter (PS) formalism [306], a Gaussian distribution of the density perturbation is assumed and is generally applicable to spherically symmetric models. In other words, the formalism depends on the shape of the perturbation profile. A physically more robust procedure to estimate the PBHs abundance has been developed known as the compaction function approach, which we here integrate in our study. The compaction function is similar to the Schwarzschild gravitational potential, since it relies mainly on the functional form of the gravitational potential(or mass excess) within a region. In [307], it was shown that the threshold of the PBHs formation is largely unaffected beyond the critical radius $r_{p}$, denoting the peak of the compaction function. A crucial takeaway of this approach is there exists a universal threshold value of the volume-averaged compaction function that can easily be related to $\delta_{c}$, and is independent of the perturbation profile to a high-accuracy estimate [307], unlike the PS method which relies on direct $\delta_{c}$ value and is heavily dependent on the perturbation shape. Additionally, PBH formation is a nonlinear process and simple consideration of a Gaussian profile of density perturbation as in PS theory is unrealistic. In contrast, the compaction function utilizes a non-linear relationship between the density and curvature perturbation in the long-wavelength approximation, which we discuss later in detail.
@@
Another important aspect of this work comes from considering the impact of quantum loop effects while studying PBH formation. More precisely, the scenario where huge enhancements in the curvature perturbation can induce significant quantum corrections at large scales whilst contributing to PBH formation was first investigated by Kristiano and Yokoyama in [308, 309]. This scenario quickly became a topic of intense discussion, followed by works such as [241, 242, 275, 274, 310, 311, 312, 313, 314, 315, 276, 273] discussing the proper conditions to conclude whether or not PBH formation remains ruled out. Parameters that include the duration of the perturbation amplifying phase, also known as the ultra-slow roll (USR) phase, and the nature (smooth or sharp) and implications of the transition in and out of this phase demand much scrutiny before establishing any firm results. In the midst of this, a series of papers by the authors of [244, 246, 245], proposed a strong no-go theorem as a constraint on the largest mass of the PBHs, $M_{\rm PBH}\sim{\cal O}({10^{2}}\rm gm)$, as a result of the USR phase. The calculation mechanism involved the application of regularization, then renormalization, followed by a dynamical group resummation, to arrive at a finite result for the two-point correlation function of importance to us in the late time. Moreover, the underlying effective field theory (EFT) treatment of the no-go theorem propelled these conclusions to cover a wide range of canonical and non-canonical single-field inflationary models.
@@
Nevertheless, several studies like [247, 248, 21, 22, 19, 20, 250, 253] have demonstrated methods to circumvent this rigid PBH mass constraint using the EFT framework. Among such diverse set-ups, our paper focuses on a specific framework with additional pre-inflationary phases of contraction and an expanding non-singular bounce before following up with the standard SR-USR-SR stages for PBH generation. For our purposes, we utilize the result of performing the same calculation scheme as explained above to obtain the regularized-renormalized-resummed (RRR) scalar power spectrum, [253], within the EFT setting. The motivation here to introduce such phases in our set-up stems mainly from the fact that other than allowing us to evade the no-go, such phases have been shown to provide non-Gaussianities (NGs) that are orders of magnitude of larger than what is attainable in a standard single-field slow-roll inflation, as first shown from Maldacena’s calculations in [316] with $f_{\rm NL}\sim{\cal O}(0.1)$, or even, in scenarios with slow-roll violating phases like the USR with $101<|f_{\rm NL}|<$ 10. The matter bounce is such a scenario where the NGs are much larger (and negative) compared to standard inflationary scenarios, primarily because the curvature perturbation here grows on scales beyond the Hubble scale which later contributes to the strength and shape of the NGs [317]. Likewise, for models focusing on ekpyrotic contraction [318], negative NGs are also shown to be larger than the standard SR inflation. At this point, it might be tempting to ask that, with the introduction of all these phases, how large of a non-Gaussianity can be realistically feasible? Before answering that question, let us briefly consider the importance of NGs in the context of PBH formation.
@@
The abundance of PBHs is affected by the Probability Distribution (PDF) of the curvature perturbation, which is generally assumed to be Gaussian. However, for a realistic analysis, the inclusion of NGs in the curvature perturbation, $\zeta$, is necessary to accurately estimate the abundance of PBHs, largely because there is no concrete confirmation that the curvature perturbation itself behaves as a Gaussian random field. They emerge from quantum fluctuations and are responsible for creating a myriad of exotic and large-scale structures in the universe. Therefore, it is highly likely that the perturbation field would inherit intrinsic NGs. Besides, even under an ideal Gaussian assumption, the relationship between the curvature perturbation and the density contrast field is non-linear, which imparts NGs to the density contrast field as an effect due to these non-linear (NL) corrections. Previous case studies in [23, 20, 19, 18, 319] demonstrate that incorporating NGs have enabled a reduction in the abundance of PBHs, eliminating the risk of overproduction. Most fiducial cases in the literature treat NGs perturbatively, while there have been some developments in its study through a non-perturbative formalism [320, 321, 322].
@@
In this paper, we will consider the local form of NGs in the curvature perturbation expressed with the parameter $f_{\rm NL}$ that encodes the first-order NG correction. PBH abundance is highly dependent on the signature of this $f_{\rm NL}$, and most cases show that $f_{\rm NL}<0$ is considered more favourable to suppress the PBH abundance while positive NGs $f_{\rm NL}>0$ invite the risk of overproduction. In our case, such NGs are affected by the USR phase, usually short-lived with $\Delta N_{\rm USR}\sim 2$, and operate to enhance the fluctuations at small scales and violating Gaussianity. One could also envision a curvaton scenario [23, 323, 324, 322], where the curvature perturbation acquires NG corrections. Apart from this, scenarios described a while ago with ekpyrotic contraction and matter bounce also invite a large amount of negative non-Gaussianity. The analysis in this paper incorporates the above-mentioned two phases along with the three phases of the usual SR inflation model. Let us revert back to the question we asked a while ago about how large of a negative NG can we dive into. In [318], it has been shown through the calculation using the $\delta N$ formalism [325, 326, 327, 328, 329, 330, 331, 332, 333, 334, 335, 336, 337, 338, 339, 340, 341, 342, 320, 343, 344, 345], that $f_{\rm NL}$ can take values as large as $f_{\rm NL}\sim-166$ from the ekpyrotic phase. While in [317], it has been shown that $f_{\rm NL}$ is negative and of the order of $f_{\rm NL}\sim{\cal O}(1)$. When these phases, along with an added USR phase, are included, a large NG is expected. For our analysis, we adopt two benchmark values of $f_{\rm NL}\in(-39.95,-35/8)$ and present a comparative analysis to conclude with a stringent bound for the leading-order NG parameter $f_{\rm NL}$. We will demonstrate through our analysis that diving below $f_{\rm NL}=-39.95$ for a specific threshold of the compaction function increases the risk of breaking perturbativity conditions. We will confirm that with $f_{\rm NL}=-39.95$, a substantial amplification of the power spectrum is possible that explains the PTA and simultaneously avoids overproduction. Throughout this work, we have employed two distinct values of the effective sound speed parameter $c_{s}\in(0.88,1)$, which lie within the window that preserves the causality and unitarity conditions. The mass spectrum of PBHs examined in this work is between $M_{\rm PBH}\sim(10^{-6}-0.1)M_{\odot}$. Our set-up is successful in generating large mass PBHs while satisfying the inflationary conditions and underlying perturbativity conditions. We also clarify that in this work the equation of state parameter has been fixed to a constant value corresponding to the radiation-dominated epoch.
@@
This paper is structured as follows. We commence with a quick description of our EFT set-up in sec.(II). This is followed by the expressions for the tree-level and loop-corrected scalar power spectrum of the curvature perturbation in sec.(III). In sec.(IV), the overall dynamics of PBHs formation has been discussed. Special emphasis has been put towards the development of the compaction function approach as it is going to be our prime tool throughout this paper. The formulas for the corresponding joint PDF and PBHs abundance have been laid out along with the domain of integration for the PBHs mass fraction. This is followed by highlighting the main expressions for the SIGW spectrum in sec.(V) for $w=1/3$. We then present the outcomes of our numerical analysis in sec.(VI), where motivation for consideration of large negative NGs for the analysis has been largely discussed. To familiarize the reader with a rough idea of the recent studies, we have presented a light comparison in sec.(VII) between several models. Various experiments that our analysis has confronted have been enlisted in sec.(VIII). Based on our examination, we provide a stringent upper bound on the NG parameter $f_{\rm NL}$, with its validity, in sec.(IX). Finally, we draw our conclusion and highlight the major findings from this work in sec.(X).